Joe is x years old.
Sue is 3 years old younger than Joe.
Paul is twice old as joe.
The sum of their ages is 33.
How old is Joe?

Found 3 solutions by saw, josgarithmetic, Alan3354:
Answer by saw(34)

Solution.
Let Joe's age be x.
Sue's age be x-3.
Paul's age be 2x.
x+(x-3)+2x = 33
Remove parentheses
x+x-3+2x = 33
Collect the like terms
4x-3 = 33
Move constant to the right
4x = 33+3
Add the numbers
4x = 36
Divide both sides of the equation by 4
x = 9
Joe is 9 years old.
Sue is (9-3) = 6 years old.
Paul is 2x = 2*9 = 18 years old.
Check-----
x+(x-3)+2x = 33

Answer by josgarithmetic(39617)

PERSON        AGE
Joe            x
Sue            x-3
Paul          2x
SUM           33

What to do is now clear enough.

Answer by Alan3354(69443)

The fact that so many people don't know how old they breaks my heart.